# NestedFP: High-Performance, Memory-Efficient Dual-Precision Floating Point Support for LLMs

### v1

@@ Abstract
Large Language Models (LLMs) are playing a crucial role in latency-critical, high-throughput services like virtual assistants and code generation. While techniques such as continuous batching and paged attention address service-level objectives (SLOs), and quantization methods accelerate inference, the dynamic and efficient adaptation of precision at runtime remains a significant, largely underexplored challenge. The emergence of hardware support for FP8 arithmetic, offering up to 2$\times$ the throughput of FP16, presents an attractive opportunity for interactive LLM serving. However, current approaches like co-deploying FP8 and FP16 models suffer from increased storage overhead and fail to unlock FP8’s full potential. To address these limitations, we introduce NestedFP, a novel precision-adaptive serving technique enabling seamless FP8 and FP16 inference from a single 16-bit model representation, thereby incurring no additional memory cost. NestedFP decomposes each FP16 weight into two 8-bit components, facilitating efficient FP8 execution while preserving full FP16 accuracy. We demonstrate the practical viability of our approach by implementing a custom CUTLASS-based GEMM kernel that reconstructs FP16 operands on-the-fly, integrated within the vLLM serving framework. Our evaluation shows that NestedFP delivers up to 1.55$\times$ throughput improvement in FP8 mode with negligible accuracy degradation compared to FP16 precision, while introducing only 3.9% performance overhead on average in FP16 mode across various models. NestedFP thus provides a flexible foundation for dynamic, SLO-aware precision selection, paving the way for more scalable and efficient LLM serving under bursty and heterogeneous workloads.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information. Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand. For example, OpenAI reportedly generates over a billion words per day—equivalent to more than one million words per second marlin —with production query rates frequently surpassing 200 requests per second scale_llm . This widespread deployment in latency-sensitive, user-facing services has intensified the need for inference infrastructures that are not only scalable and efficient, but also capable of satisfying stringent latency and throughput constraints.
The latency of LLM inference is commonly evaluated using two standard metrics: Time To First Token (TTFT), which measures the prefill latency, and Time Per Output Token (TPOT), which captures the average latency for generating each output token. For interactive applications, industry-standard service-level objectives (SLOs) typically require TTFT to be under 200 ms and TPOT to remain below 33.3 ms in order to ensure a responsive user experience baseten_slo . To meet these stringent latency targets under growing demand, several optimization techniques have been proposed. ORCA orca introduces efficient batching mechanisms that consolidate linear layer execution across both the prefill and decode phases. PagedAttention paged_attention addresses memory inefficiencies arising from unpredictable output sequence lengths by reducing over-allocation, thereby increasing the number of concurrently served requests. More recently, Sarathi-Serve sarathi_serve proposes chunked execution of the prefill phase and constructs hybrid batches combining prefill and decode requests to further improve scheduling efficiency.
Recent studies have shown that the FP8 format can achieve accuracy comparable to FP16 across a variety of large language model tasks fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 . The introduction of native FP8 support on modern accelerators—such as NVIDIA Hopper GPUs and Intel Gaudi HPUs—has made FP8 increasingly practical for large-scale inference. On NVIDIA Hopper, FP8-compatible tensor cores can offer up to 2× higher peak throughput compared to their FP16 counterparts, making FP8 particularly advantageous for latency-sensitive or bursty workloads. In this work, we propose leveraging both FP8 and FP16 representations to better satisfy service-level objectives (SLOs). However, co-deploying models in both formats is non-trivial, as it is difficult to ensure robust accuracy and performance without introducing additional storage overhead beyond the originally required 16-bit model.
@@
We present NestedFP, a technique that enables dynamic selection between FP8 and FP16 precision from a single, unified model representation. By analyzing the value distribution of FP16 weights, our approach restructures each weight tensor into two 8-bit components, enabling accurate FP16 reconstruction while supporting efficient FP8 execution—all with no additional memory storage. To realize this format, we develop a custom General Matrix Multiplication (GEMM) kernel built on the CUTLASS library cutlass , which reconstructs FP16 values from the two 8-bit components on-the-fly during GEMM kernel execution. The kernel incurs an average performance overhead of 6.47% across varying GEMM shapes, relative to the vanilla CUTLASS kernels, respectively. Moreover, our 8-bit tensor for FP8 inference achieves FP8 inference accuracy comparable to that of standard FP8 quantized models.
We integrate our custom kernel into the vLLM serving framework vllm and show that it incurs only 3.90% end-to-end performance overhead on average across models in FP16 mode, while delivering up to 1.55$\times$ performance improvements when executing in FP8 mode. NestedFP provides a flexible and memory-efficient foundation for adaptive precision serving, making it well-suited for fluctuating request rates and stringent latency constraints.
@@ 2.1 Floating Point Representation
[…]
@@ 2.2 FP8 Quantization for LLMs
Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs investigation_fp8_across_accelerators . Leveraging this support, FP8 GEMM operations can achieve up to 2×2$\times$2 × speedup compared to their FP16 counterparts. This growing availability is accelerating the adoption of floating-point quantization—particularly FP8—further highlighting its relevance for efficient LLM deployment.
@@
FP8 Results. While FP8 quantization offers significant efficiency benefits and shows promising results, it still lags behind FP16 in terms of accuracy. Table 1 presents a comparison between FP8 and FP16 across various benchmarks that require sophisticated reasoning capabilities. The FP8 results are obtained using the vLLM framework vllm , with E4M3 quantization applied using per-channel weight scaling and per-token activation scaling based on absolute maximum value. Quantization is applied exclusively to linear layers, with embedding layers left in higher precision. FP8 generally exhibits a measurable drop in accuracy, highlighting its current limitations for tasks that demand high output fidelity.
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ Minerva Math 18.1 17.6 --0.54 17.1 16.9 --0.20 42.7 42.9 $+$0.16 35.4 34.6 --0.84 MMLU Pro 33.3 32.8 --0.49 35.4 34.9 --0.44 53.1 52.7 --0.37 54.3 53.6 --0.71 BBH 39.5 38.6 --0.83 41.7 41.2 --0.49 27.5 27.1 --0.38 52.5 51.8 --0.65
@@ 3 Motivation
3.1 Challenge: Dynamic Resource Demand in LLM Serving
LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests andes ; hygen . These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in TTFT and TPOT. Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet service-level objectives (SLOs).
To quantify the extent of this variability, we analyzed production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace . These traces contain timestamps of incoming requests, from which we constructed the per-second request rate on May 10, 2024, as shown in Figure 1. Our analysis reveals that the request rate ranged from 0 to 100 requests/s across the entire day, and exhibited intra-day variations of up to 5.8-fold within the most variable 1-hour window (14:00–15:00 UTC; min: 17, max: 98) and 3.2-fold within the most variable 1-minute window (18:12–18:13 UTC; min: 31, max: 98).
To cope with such highly variable request patterns, auto scaling is a commonly used to automatically adjusts computational resources based on real-time demand auto_scaling . However, auto scaling is feasible mainly in cloud environments with sufficient available resources, and is typically limited to time scales of few minutes to hours auto_scaling_timescale . Therefore, it struggles to respond effectively to sharp second-level spikes and is infeasible for adjusting hardware at millisecond-level intervals.
In contrast, LLM serving systems operate on a continuous batching basis orca , where requests are scheduled at the granularity of iteration steps usually every tens to hundreds of milliseconds. As the batch size of each step directly impacts token-level latency, variability in request rates must therefore be evaluated at the same short-term timescale. Therefore, high variability within second-level time scales can cause substantial fluctuations in TTFT and TPOT, underscoring the need for serving strategies that can operate effectively at finer timescales, independent of slower hardware scaling mechanisms.
Figure 1: (a) Request rates fluctuations observed on May 10, 2024 from Azure LLM inference trace. (b) Comparison of p90 TPOT across different model precisions, using the Azure LLM inference trace recorded on May 10, 2024.
@@ 3.2 Our Proposal: Dual-Precision LLM Serving
To address short-term fluctuations in resource demand during LLM serving, we propose a simple yet effective dual-precision strategy that dynamically switches between FP16 and FP8 execution. When system load is low, FP16 is used to preserve maximum model quality. Under high load conditions, the system falls back to FP8 execution to reduce computational and memory overhead, thereby maintaining responsiveness. Although FP8 may incur slight accuracy degradation, it is preferable to violating SLOs during periods of request surges.
Figure 1 shows p90 TPOT under three precision modes—FP16, FP8, and our dual-precision scheme—using the Llama 3.1 8B model with vLLM on an H100 GPU. Request patterns are based on Microsoft Azure’s LLM inference trace azure_llm_inference_trace , downscaled to 20%. During the 60-second interval with fluctuating load (1–11 req/s, avg. 5.05 req/s), FP16 incurs 19 seconds of SLO violation (TPOT > 33 ms), while FP8 reduces this to 8 seconds. Our dual-precision strategy maintains the same SLO compliance as FP8 while preserving FP16 accuracy for over 68% of the time. This demonstrates its effectiveness in improving latency robustness without sacrificing model quality.
@@ 3.3 How to Implement Dual-Precision LLM Serving?
To enable adaptive use of both FP8 and FP16 precision in LLM serving, two straightforward implementation strategies can be considered: separate storage and on-the-fly dequantization.
Separate Storage. One simple approach is to maintain both FP8 and FP16 models in memory and switch between them based on task requirements. While this enables the on-demand selection between lower- and higher-precision models, it results in a 50% memory overhead—an impractical burden for LLM serving systems already constrained by model size. This memory capacity overhead not only limits the maximum model size that can be deployed but also reduces the available memory for key-value (KV) caching, thereby restricting the number of concurrent requests and degrading overall serving throughput.
On-the-fly Quantization. An alternative approach is to store only the FP16 model and perform on-the-fly quantization to enable execution on FP8 tensor cores. Unlike the separate storage method, this approach avoids additional memory capacity overhead. However, it can lead to severely suboptimal GEMM performance in FP8 mode. Despite executing in 8-bit precision, the full 16-bit weights must still be fetched from memory, effectively doubling memory traffic compared to native FP8 execution. As a result, memory bandwidth becomes a bottleneck, limiting the potential throughput gains of FP8 and negating its advantages over the FP16 baseline.
These limitations highlight the need for a more sophisticated dual-precision serving mechanism—one that eliminates additional memory overhead while enabling efficient GEMM execution in both FP16 and FP8 modes.
@@ 4 NestedFP
4.1 NestedFP Overview
For efficient dual-precision LLM serving, we propose NestedFP, a scheme that stores only 16-bit weights in memory and extracts 8-bit upper tensor from them when operating in FP8 mode. This approach not only avoids additional memory usage, but also prevents memory traffic amplification—such as fetching 16-bit weights when performing 8-bit GEMM. Figure 2 illustrates the core concept of NestedFP, which addresses the following two key challenges to make this approach viable.
First, the extracted 8-bit representation must maintain accuracy comparable to existing FP8 quantization methods. Naively truncating the upper 8 bits of FP16 values (which use the E5M10 format) produces a low-precision representation similar to E5M2, which offers limited precision compared to the commonly preferred E4M3 format for LLM inference. Note that FP16 is basically E5M10. Moreover, this naive truncation often performs worse due to indiscriminate rounding, further degrading accuracy. This motivates the need for a quantization-aware format that simultaneously supports both FP16 and FP8 execution. Section 4.2 elaborates on this data format design.
Second, kernel efficiency must be maintained across both precisions. Achieving this requires storing the two 8-bit halves of each 16-bit weight as separate tensors. On H100 GPUs, where DRAM is accessed in 64-byte units, extracting individual bytes from packed 16-bit data results in wasted bandwidth. Separating the 8-bit components enables efficient FP8 execution using a single tensor, while FP16 models require reconstructing the original 16-bit weights via a custom GEMM kernel that dynamically fuses the components at runtime. Section 4.3 details this kernel design.
Figure 2: FP16 weight is first pre-processed into 8-bit Upper and Lower parts. Both parts are used for FP16 model execution while only Upper part is used for FP8 model execution.
@@ 4.2 Compact yet Effective Data Format for Dual-Precision LLM
Opportunities: Low-Entropy Exponent Bit in FP16 Models. Figure 3 shows the the weight distributions of FP16 linear layers—including QKVO projections and MLPs—across four different LLMs. The vast majority of weight values have absolute magnitudes less than or equal to 1.75, which means that their most significant exponent bits in FP16 are zero. Specifically, as shown in Figure 3, three of the four evaluated models exhibit this trait across all layers. The sole exception, Phi-4, contains some layers that deviate from this pattern, but these account for only 8.75% of its layers. This strong tendency suggests an opportunity for lossless exponent mapping from FP16 to the FP8 E4M3 format, which also uses 4 exponent bits.
Min Max NestedFP Applicable Layers
Llama 3.1 (8B) -0.92 0.91 128 of 128 (100%)
Mistral Nemo Base (12B) -0.92 0.95 160 of 160 (100%)
Phi-4 (14B) -3.60 3.89 146 of 160 (91.25%)
Mistral Small Base (24B) -0.49 0.45 160 of 160 (100%)
@@
Figure 3: (a) Weight distributions of linear layers across different models. (b) Per-model weight range (min/max) and count of linear layers eligible for NestedFP quantization.
Our Proposal. Leveraging the unused exponent bit in FP16, NestedFP restructures FP16 values so that high-quality E4M3 representations can be directly extracted. Figure 4a illustrates the refactoring process. The lower 8 bits of new format of NestedFP is simply constructed by inheriting lower 8 bits of the FP16 mantissa. The upper 8 bits are formed from the sign bit and 4 exponent bits (excluding the most significant bit of the FP16 exponent) of the original FP16 value, along with a 3-bit mantissa obtained by quantizing its 10-bit mantissa.
This simple mapping scheme introduces a fixed scaling factor determined by the difference in exponent bias between the FP16 and E4M3 formats, which is $2^{8}$. In other words, NestedFP effectively applies E4M3 quantization with a scaling factor of $2^{8}$. To quantize the 10-bit FP16 mantissa down to 3 bits, we employ a round-to-nearest-even strategy. Specifically, we examine the 7 least significant bits of the FP16 mantissa (i.e., bits M[4:10]) to determine whether the value lies above or below the midpoint (64). If the value is greater than 64, we round up. If the value equals 64, we round up only when the least significant bit of the resulting 3-bit mantissa (M3) is 1. If the value is less than 64, we round down. This division process is done offline prior to model execution.
@@
FP8 GEMM can be performed directly using the upper 8 bits, whereas FP16 GEMM requires online reconstruction of 16-bit values, introducing additional complexity. Figure 4b illustrates this reconstruction process. Essentially, it concatenates the upper 8 bits and lower 8 bits, but not by simple concatenation. The earlier rounding decision must be reversed by checking an implicit checksum; if the LSB of the upper 8 bits differs from the MSB of the lower 8 bits, a rounding-down correction is applied before recombining the exponent and mantissa bits into a valid FP16 value. This process ensures precise, lossless recovery of the original FP16 values.
@@
Figure 4: (a) Pre-processing FP16 weight value into two 8-bit NestedFP components (Offline). (b) On-the-fly reconstruction of FP16 weight values for tensor core execution (Online).
Handling Exception Layers. Since most models do not contain weights with absolute values exceeding 1.75, we leave these values in their original FP16 format without pre-processing. As a result, the corresponding layers are always executed in FP16, consistent with standard FP16 model execution. As shown in Sections 5, this exception handling preserves both model accuracy and end-to-end performance, as these layers constitute only a small portion of the total linear layers.
4.3 GPU Kernel Design

### v2

@@ Abstract
Meeting service-level objectives (SLOs) in Large Language Models (LLMs) serving is critical, but managing the high variability in load presents a significant challenge. Recent advancements in FP8 inference, backed by native hardware support, offer a potential solution: executing FP16 models by default, while switching to FP8 models during sudden load surges to achieve higher throughput at the cost of a slight quality degradation. Although this approach facilitates effective SLO management, it introduces additional memory overhead due to storing two versions of the same model. In response, this paper proposes NestedFP, an LLM serving technique that supports both FP16 and FP8 models in a memory-efficient manner by overlaying FP8 parameters onto FP16 parameters, allowing both models to share the same FP16 memory footprint. By leveraging a compact data format for the overlay and a specialized GEMM kernel optimized for this format, NestedFP ensures minimal degradation in both model quality and inference throughput across both FP8 and FP16 modes. NestedFP provides a flexible platform for dynamic, SLO-aware precision selection. The code is available at https://github.com/SNU-ARC/NestedFP.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information gpt4_technical_report ; gpt3_technical_report ; llama3_technical_report . Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand. For example, OpenAI’s production-scale models process roughly 100 billion words per day openai_billion_token and production query rates frequently surpass 200 requests per second chatgpt_fact ; scale_llm .
To accommodate such large-scale applications, research efforts have focused not only on improving serving throughput orca ; paged_attention but also on meeting service-level objectives (SLOs) sarathi_serve ; llumnix ; sola_llm . One of the key challenges in achieving SLOs for LLM serving lies in handling dynamic load. Request rates, as well as input and output sequence lengths, can fluctuate significantly over short time intervals. Simply overprovisioning compute resources may mitigate this issue but results in substantial system cost inefficiency.
As a promising alternative, we explore an approach we call dual-precision LLM serving, which utilizes FP8 together with FP16, the de facto standard number format for weight parameters. While FP8 may incur a slight degradation in model quality, it offers up to 2× higher peak throughput compared to FP16 fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 . By operating in FP16 mode under normal conditions and switching to FP8 mode only during sudden load surges, we can effectively manage dynamic workloads without resource overprovisioning. This approach achieves a better balance between SLO attainment and service quality than relying solely on either FP16 or FP8.
Co-deploying models in both formats, however, is non-trivial. Simply keeping both in memory is infeasible due to excessive memory capacity overhead. Storing only FP16 weights and quantizing them to FP8 on the fly when needed results in highly suboptimal FP8 throughput. Therefore, an effective solution is needed to support dual-precision LLM serving without incurring memory overhead or throughput degradation.
@@
In this paper, we present NestedFP, a technique that enables efficient dynamic selection between FP8 and FP16 precision from a single unified model representation. NestedFP decomposes each 16-bit weight tensor into two 8-bit components, allowing FP8 inference without additional memory overhead or throughput degradation. Since FP16 mode requires reconstructing 16-bit values by merging the two 8-bit components in this approach, we develop a custom General Matrix Multiplication (GEMM) kernel, built on the CUTLASS library cutlass , that performs accurate on-the-fly FP16 reconstruction during kernel execution. According to our evaluation, NestedFP enables high-quality, dual-precision LLM serving efficiently: FP8 models generated from decomposed FP16 weights match the quality of existing FP8 quantization methods, and our FP16 GEMM kernel incurs only a 6.47% average performance gap relative to the CUTLASS baseline, which translates to 4.98% end-to-end overhead across models.
Our contributions are summarized as follows:
  • We propose dual-precision LLM serving as an effective strategy to handle dynamic load fluctuations.
  • We introduce a novel data format that supports both high-quality FP16 and FP8 inference in a memory-efficient manner.
  • We develop a custom FP16 GEMM kernel for this data format, which performs on-the-fly FP16 value reconstruction during execution.
  • We demonstrate that NestedFP enables high-quality dual-precision LLM serving with no memory overhead and high throughput, significantly improving SLO attainment compared to FP16-only baselines.
@@ 2.1 Floating Point Representation
A floating point format with $x$ exponent bits and $y$ mantissa bits is denoted as ExxMyy. The exponent width $x$ determines the dynamic range of representable values, while the mantissa width $y$ governs numerical precision. A floating point value $X_{FP}$ in this format can be expressed as shown in Equation 1, where $S$ is the sign bit, $M_{j}$ are the mantissa bits, $E$ is the unsigned integer formed by the exponent bits, and $b$ is the exponent bias.
@@ 2.2 FP8 Quantization for LLMs
Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs investigation_fp8_across_accelerators . Leveraging this support, FP8 GEMM operations can achieve up to $2\times$ speedup compared to their FP16 counterparts. This growing availability is accelerating the adoption of floating-point quantization—particularly FP8—further highlighting its relevance for efficient LLM deployment.
@@
FP8 Results. While FP8 inference offers significant throughput gains, it comes with the potential caveat of quality degradation due to the inevitable loss of information associated with reduced precision dfloat11 . Table 1 presents a comparison between FP8 and FP16 across various models and benchmarks. Specifically, for FP8 quantization, we use the E4M3 format with per-token activation and per-channel weight quantization for all linear layers, using the vLLM framework vllm . A consistent accuracy degradation is observed in FP8 models, which clearly shows the trade-off between throughput and model quality in FP8 quantization.
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ Minerva Math 18.2 17.6 -0.60 17.2 16.5 -0.72 42.7 42.9 +0.14 35.4 34.2 -1.14 MMLU Pro 33.3 32.8 -0.49 35.4 35.3 -0.06 53.1 52.7 -0.36 54.3 53.6 -0.77 BBH 39.5 38.6 -0.86 41.6 40.6 -0.97 27.5 27.1 -0.40 52.4 50.8 -1.62
@@ 3 Motivation
3.1 Challenge: Dynamic Load in LLM Serving
LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests andes ; hygen . These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in Time To First Token (TTFT) and Time Per Output Token (TPOT). Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet SLOs.
To quantify the extent of this variability, we analyze production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace . Specifically, we measure how the request rate fluctuates over a one-hour interval (00:00–01:00 UTC, May 10, 2024, in the trace). Figure 1 presents the results. The request rate exhibits nearly a fivefold variation between the lowest and highest load periods. Moreover, as shown in the one-minute zoom-in view, these fluctuations occur at high temporal frequency, indicating second-level variability.
A classic approach to handling such load variations is auto-scaling, which adaptively adjusts computational resources (e.g., increasing or decreasing the number of GPUs in a cluster) auto_scaling . However, auto-scaling is primarily effective in cloud environments with abundant resources and typically operates on minute-to-hour timescales auto_scaling_timescale . Consequently, it is inadequate for LLM serving systems, which must respond to sharp, second-level load spikes and scale throughput within millisecond-level intervals to meet SLO requirements.
Figure 1: (a) Request rate fluctuations during the first hour of May 10, 2024, from the Azure LLM inference trace. The inset plot highlights 1-minute variability from the 30-minute mark. (b) Comparison of p90 TPOT across different model precisions, using the Azure LLM inference trace recorded on May 10, 2024.
@@ 3.2 Our Proposal: Dual-Precision LLM Serving
To address short-term fluctuations in resource demand during LLM serving, we propose a simple yet effective dual-precision strategy that dynamically switches between FP16 and FP8 modes. When the system load is low, FP16 is used to preserve maximum model quality. Under high load conditions, the system falls back to FP8 mode to increase throughput, thereby maintaining responsiveness. Although FP8 mode may incur slight accuracy degradation, it is often preferable to violating SLOs during periods of request surges.
Figure 1 presents the per-second p90 TPOT under three precision schemes: FP16, FP8, and the proposed dual-precision format, using the Llama 3.1 8B model on an H100 GPU with vLLM. The request patterns are extracted from a 60-second segment of Microsoft Azure’s LLM inference trace azure_llm_inference_trace , scaled down to 20% of the original load. The TPOT SLO threshold is set to 33 ms baseten_slo . As shown, the FP16 model exhibits more frequent TPOT spikes that exceed the TPOT SLO threshold compared to the FP8 model. To further quantify the difference, we analyze the rate of SLO violations across the same trace: 35.8% of requests under FP16 exceeded the threshold, whereas FP8 reduced this to 17.3%. In short, the proposed dual-precision inference can achieve FP8-level SLO compliance while minimizing quality degradation.
@@ 3.3 How to Implement Dual-Precision LLM Serving?
To enable dynamic switching between FP16 and FP8 modes in LLM serving, two straightforward implementation strategies can be considered: separate storage and on-the-fly dequantization.
Separate Storage. One simple approach is to maintain both FP16 and FP8 models in memory and switch between them according to system load fluctuations. While this enables fast switching, it introduces a 50% memory overhead—a significant burden for LLM serving systems already constrained by model size. This memory capacity overhead not only limits the maximum deployable model size but also reduces the available memory for key-value (KV) caching, thereby restricting the number of concurrent requests and ultimately degrading overall serving throughput.
On-the-fly Quantization. An alternative approach is to store only the FP16 model and perform on-the-fly quantization to perform computation in FP8. Unlike the separate storage strategy, this approach avoids additional memory capacity overhead. However, its FP8 performance is suboptimal; although computation is carried out by FP8 units with high throughput, the full 16-bit weights must still be loaded from memory, effectively doubling memory traffic compared to native FP8 execution. As a result, memory bandwidth becomes a bottleneck, limiting the achievable throughput and diminishing much of FP8’s advantage over the FP16 baseline.
These limitations highlight the need for a more sophisticated dual-precision serving mechanism—one that eliminates additional memory overhead while enabling efficient GEMM execution in both FP8 and FP16 modes.
@@ 4 NestedFP
Overview. Figure 2 illustrates the core concept of NestedFP, which stores only 16-bit weights in memory and extracts 8-bit weights from them when operating in FP8 mode. Specifically, NestedFP decomposes each FP16 parameter into two parts: the upper 8 bits and the lower 8 bits. With these two parts of the weight parameters stored as separate tensors, both parts are loaded in FP16 mode, whereas only the upper part is loaded in FP8 mode. This approach not only eliminates additional memory usage but also prevents memory traffic amplification—for example, fetching 16-bit weights when executing in 8-bit mode.
Challenge 1: 8-bit Mode Accuracy. One key challenge of this approach is ensuring that the 8-bit representation extracted from 16-bit parameters achieves sufficient accuracy. In other words, it must maintain performance comparable to existing FP8 quantization methods. Naive truncation (simply using the upper 8 bits of FP16 values) yields a representation similar to E5M2, which offers limited precision compared to the commonly preferred E4M3 format for LLM inference. Moreover, such truncation often performs worse due to indiscriminate rounding, further degrading accuracy. This motivates the need for a quantization-aware data format that enables the extraction of high-quality FP8 representations. Section 4.1 elaborates on this data format design.
Challenge 2: 16-bit Mode Throughput. Another key challenge is maintaining throughput in the 16-bit mode. The 8-bit mode in NestedFP trivially achieves high performance, as native FP8 GEMM kernels can be used directly after loading the upper-part tensor. However, the 16-bit mode is less straightforward: both the upper and lower-part tensors must be loaded and combined to reconstruct the original 16-bit weights. This requires a custom GEMM kernel capable of performing the reconstruction dynamically. Section 4.2 details this kernel design.
Figure 2: FP16 weight is decomposed into 8-bit upper and lower parts. Both parts are used for FP16 mode while only upper part is used for FP8 mode.
@@ 4.1 Compact yet Effective Data Format for Dual-Precision LLM
Opportunities: Low-Entropy Exponent Bit in FP16 Models. Figure 3 shows the the weight distributions of all FP16 linear layers across four different LLMs. The vast majority of weight values have absolute magnitudes less than or equal to 1.75, which means that their most significant exponent bits in FP16 are zero. Specifically, as shown in Figure 3, three of the four evaluated models exhibit this trait across all layers. The sole exception, Phi-4, contains some layers that deviate from this pattern, but these account for only 8.75% of its layers. This strong tendency suggests an opportunity for lossless exponent mapping from FP16 to the FP8 E4M3 format, which also uses 4 exponent bits.
Min Max NestedFP Applicable Layers Llama 3.1 (8B) -0.92 0.91 128 of 128 (100%) Mistral Nemo Base (12B) -0.92 0.95 160 of 160 (100%) Phi-4 (14B) -3.60 3.89 146 of 160 (91.25%) Mistral Small Base (24B) -0.49 0.45 160 of 160 (100%)
@@
Figure 3: (a) Weight distributions of linear layers across different models. (b) Per-model weight range (min/max) and the proportion of linear layers where NestedFP can be applied.
Our Proposal. Figure 4 illustrates how NestedFP decomposes an FP16 weight such that the upper 8 bits can serve as a high-quality E4M3 representation by leveraging an unused exponent bit in FP16. First, the lower 8 bits directly inherit the lower 8 bits of the FP16 mantissa. The upper 8 bits, which correspond to the E4M3 representation, are constructed as follows. The sign bit is directly inherited from FP16. The next four exponent bits are also copied from FP16, excluding the most significant exponent bit, which is assumed to be unused (or zero). Finally, 3 most significant mantissa (i.e., bits M[1:3]) bits are obtained and constructed value is applied rounding using a round-to-nearest-even policy. Specifically, the 7 least significant bits of the FP16 mantissa (i.e., bits M[4:10]) are examined to determine whether the value lies above or below the midpoint (64). If the value is greater than 64, it is rounded up; if it equals 64, it is rounded up only when the least significant bit of the resulting 3-bit mantissa (M3) is 1; otherwise, it is rounded down. This decomposition process is performed offline prior to model execution.
In this scheme, removing most significant exponent bit effectively acts as multiplying $2^{8}$ between E4M3 and FP16 value. In other words, NestedFP effectively applies E4M3 quantization with a fixed scaling factor of $2^{8}$.
@@
FP16 Reconstruction. Figure 4 illustrates how NestedFP reconstructs FP16 weights at runtime. Essentially, this process merges the upper and lower 8-bit parts, but not through a simple concatenation. The rounding decision made during decomposition must be reversed by checking an implicit checksum; if the LSB of the upper 8 bits differs from the MSB of the lower 8 bits, a rounding-down correction is applied before recombining the exponent and mantissa bits into a valid FP16 value. This process ensures precise, lossless recovery of the original FP16 weights.
@@
Figure 4: (a) Decomposition of FP16 weight values into two 8-bit parts (offline). (b) Reconstruction of FP16 weight values (online).
Handling Exception Layers. For layers containing weight values with magnitudes exceeding 1.75 (i.e., where the MSB of the exponent is non-zero), we retain them in FP16 and always perform FP16 computation. This may reduce the benefits of dual-precision LLM serving, as the throughput advantage of FP8 mode becomes less pronounced. However, it is important to recall that, as shown in our analysis of popular models (Figure 3), such layers are rare and therefore have minimal impact on the overall effectiveness of the dual-precision scheme.
4.2 FP16 GEMM Kernel with On-the-fly Reconstruction